OpenStudy (anonymous):

What is an equation of the line in slope-intercept form? m=2/7 and the y-intercept is (0,-12).

OpenStudy (anonymous):

1.y=2/7x+12 2.y=2/7x-12 3.y=12x-2/7 4.y=12x+2/7

OpenStudy (anonymous):

y intercept = b = (0, -12) slope = m = 2/7 y = mx + b = slope intercept form y = 2/7x - 12

OpenStudy (anonymous):

Thank you so much, could you possibly help me with some more questions?